OpenStudy (lilsis76):

An 8.5g Ice Cube is placed into 255g water. Calculate temperature change in water upon the complete melting of the ice. (Assume all the energy required to melt the ice comes from the water. )

OpenStudy (aaronq):

okay for this you have to use this equation "q=m*C*\(\Delta T\)", and, q=\(\Delta H_{fusion}*n\). remember that this last one can be n or m, depending on what the units for \(\Delta H_{fusion}\) are. ------------------------------------------------------- Also, remember that \(\Delta T=T_f-T_i\) where, \(T_f\)=final temperature and \(T_i\)=initial temperature although this won't be used for this specific question ------------------------------------------------------- So, what you wanna do is to determine what you're ultimately looking for. The question asks for "Calculate temperature change in water", which is \(\Delta T\). So, first you have to find the heat (q) required to melt the ice (which will come from the water). so convert the mass of ice to grams, then mass to moles. (look up a conversion chart on google if that helps). and use \(q_{ice}=\Delta H_{fusion}*n\) (\(\Delta H_{fusion}\) needs to be looked up) Then, since all the heat to melt the ice came from the water, we express it mathematically, like this: \(q_{ice}=-q_{water}\) so, \(q_{ice}=-q_{water}=m*C*\Delta T\) Here you're using the heat capacity (C) for water, the mass of the water (not including the ice) and you're solving for \(\Delta T\).

OpenStudy (aaronq):

EDIT: sorry i had a mental relapse when i said "convert the mass of ice to grams, then mass to moles." It should only be "mass to moles".

OpenStudy (lilsis76):

haha okay im almost done reading what you put.

OpenStudy (aaronq):

okay.

OpenStudy (lilsis76):

okay im on the first step let me look in the book or heat required to melt the ice unless its 40.7 kj/mol

OpenStudy (aaronq):

no thats to vaporize (turn into steam).

OpenStudy (lilsis76):

okay so.. in the book to melt a solid ice cube to water is....2) Ice melting to liquid = 6.02 kJ/mol

OpenStudy (aaronq):

okay, so convert the mass of ice to moles

OpenStudy (lilsis76):

so i could do the 6.02 kj/mol // 1000 kj= .00602 mol

OpenStudy (lilsis76):

right?

OpenStudy (aaronq):

Just follow what i said

OpenStudy (aaronq):

n=m/M= 8.5 g /18g/mol =

OpenStudy (lilsis76):

what does the q mean again?

OpenStudy (aaronq):

heat

OpenStudy (lilsis76):

okay .472 moles

OpenStudy (aaronq):

okay, not find the heat required to melt that much ice q=\(\Delta H_{vap}*n\)

OpenStudy (aaronq):

now find*

OpenStudy (lilsis76):

okay ... .472mol (6.02 kj/ 1 mol)=2.84

OpenStudy (aaronq):

good stuff! don't forget the units 2.84 kJ now, the heat required to melt the ice comes from the water, so, \(q_{ice}=- \;q_{water}\) 2.84 kJ = -2.84 kJ use the second formula, -2.84 kJ = m*C*\(\Delta T\) you should convert 2.84 kJ to J (multiply by 1000).

OpenStudy (lilsis76):

-2840J

OpenStudy (aaronq):

nice, now solve for \(\Delta T\), -2840 J = m*C* \(\Delta T\) C=4.184 J/\(C^{\circ}g\)

OpenStudy (lilsis76):

okay @aaronq i got -5608.56 moles

OpenStudy (aaronq):

your answer should be in terms of degrees celsius because you're solving for the change in temperature, not moles.

OpenStudy (aaronq):

so you have 255 g of water -2840 J = (255 g)*(4.184 J/C*g)*\(\Delta T\)

OpenStudy (lilsis76):

sorry, celcius

OpenStudy (lilsis76):

hey @aaronq I have to go. I will continue this tomorrow. Have a goodnite!

OpenStudy (aaronq):

okay, see ya !

OpenStudy (lilsis76):

Hi @aaronq Sorry I had go go, my folks got back home, and I didnt want them to see me crying.

OpenStudy (lilsis76):

Im ready, refreshed and Im ready to go :)

OpenStudy (aaronq):

So did you get that last part?

OpenStudy (lilsis76):

yes i did :) it is wait haha

OpenStudy (lilsis76):

(-2840 J) = (255g)(4.184 J/ C*g) (delta T) I rearranged to get the delta T(temperature change) to one side so it looks like this, right? \[\Delta T = -2840 J (255g)(4.184\frac{ J }{ C*g }\]

OpenStudy (aaronq):

not quite, you're algebraically solving. A*B=C ->solving for A, A=\(\dfrac{C}{B}\) right?

OpenStudy (lilsis76):

yes, okay so then it would look like this, hold on

OpenStudy (lilsis76):

\[-2840 J / (255 g) (4.184 \frac{ J }{ C * g } )\]

OpenStudy (aaronq):

thats' right. notice that the units "g" and "J" will cancel out.

OpenStudy (lilsis76):

okay they cancel

OpenStudy (lilsis76):

i got then

OpenStudy (lilsis76):

-2.66 Celcius

OpenStudy (aaronq):

okay good stuff.

OpenStudy (lilsis76):

is this correct? then Im guessing we have to do another calculation, right? @aaronq

OpenStudy (aaronq):

no that's it, the question is asking you for the change in temperature. The change in temperature will be -2.66 celsius, since you're not given the temp of the water, you can't say what it's final temp will be.

OpenStudy (lilsis76):

oh thats true..................... YOURE SO AWESOME! :D

OpenStudy (aaronq):

haha thanks ! chemistry is my strong point

OpenStudy (lilsis76):

it really is haha okay, I have another Problem, I will close this one. thank you very mucH!

OpenStudy (aaronq):

no problem! okay

OpenStudy (lilsis76):

when I try messaging you it says only accepts people that are fanned, what does that mean? I was thinking about sending you a message if and when I need help later on.

OpenStudy (lilsis76):

@aaronq

OpenStudy (aaronq):

yeah, sorry about that. I switched my settings so only people i'm a fan of will be able to msg me because i used to get A LOT of messages of people asking for help, it got pretty annoying. But i've fanned you, now you can msg me! lol

OpenStudy (lilsis76):

okay, thank you, i was like WHA?! 0.o haha

OpenStudy (lilsis76):

WAIT!!! i just saw the rest of the quesiton to the problem haha we found the temperature change. it says calculate temperature change in water upon the complete melthing of the ice. Assume all the energy required to melt the ice comes from the water. So then that -2.66 C is still the answer right? I just want to double chekc.

OpenStudy (aaronq):

yep, we calculated the heat needed to melt the ice, which would be extracted from the liquid water, and we found the change of temperature of the water.

OpenStudy (lilsis76):

okay, haha I was gonna freak when I saw that i was like wait aminute did i forget something haha.

OpenStudy (lilsis76):

thank you Aaronq. Im going to close this problem, i have to clean up the house to give myself a quick break.

OpenStudy (aaronq):

okay! clearing up your head with some other activity is a good idea, see ya

OpenStudy (lilsis76):

thank you Aaronq. Im going to close this problem, i have to clean up the house to give myself a quick break.